# Large language models surpass human experts in predicting neuroscience results

### v2

@@ Acknowledgments
This work was supported the ESRC (ES/W007347/1), Microsoft (Accelerate Foundation Models Research Program), and a Royal Society Wolfson Fellowship (18302) to B.C.L. We thank Mona Garvert, Pradeep Reddy Raamana, Todd Hare, Yoav Kessler, and Oliver Robinson for their assistance. We thank the participants of the online study (see Table S.5).
@@
Co-authors (Table S.4) and GPT-4 (Azure OpenAI API; version 2023-05-15) created test cases that formed BrainBench. All test cases were sourced from Journal of Neuroscience abstracts published in 2023. The abstracts are organized into five sections, namely Behavioral/Cognitive, Systems/Circuits, Neurobiology of Disease, Development/Plasticity/Repair, and Cellular/Molecular. In constructing BrainBench, we incorporated a total of 200 test cases crafted by human experts and an additional 100 test cases generated by GPT-4 (Azure OpenAI API; version 2023-05-15). All test cases were subjected to extensive quality control by human experts and GPT-4. For the distribution of test cases among subfields, refer to Fig. 3 for human-created cases and Fig. S.11 for GPT-4 generated cases.
@@
We tested LLMs by adapting the Eleuther AI Language Model Evaluation Harness framework (?), which evaluates LLMs using a multiple choice setting. We presented LLMs with two versions of the abstracts from each test case separately. We prefixed each abstract with the prompt “You are a neuroscientist with deep knowledge in neuroscience. Here is an abstract from a neuroscience publication:” and applied model-specific instruction templates where appropriate. We then measured the perplexity of both passages and used perplexity as the indicator of whether LLMs favor one abstract or the other.
@@
To address this concern, we employed a common method from the literature to determine whether a given text is part of LLM’s training data (?, ?). This method involves calculating the zlib entropy and the perplexity ratio (Eq. 3) of a text sequence to infer its membership status.
@@
Zlib entropy is computed using the zlib text compression algorithm (?), which measures the level of uncertainty in a text when compressed. It’s a data-agnostic way of evaluating text. On the other hand, LLM perplexity depends on the specific training data and thus is data-dependent. In general, if a piece of text surprises zlib but not LLM, it’s likely part of the training data.
@@
RefinedWeb Known to be in Falcon’s training set (?)
@@
Arxiv (Jun-Dec 2023) Known to not be in Llama2 and Galactica’s training set (?, ?)
Biorxiv (Jun-Dec 2021) Known to be in Galactica’s training set (?)
@@
Human Evaluation
Participants
Procedure
@@
Subsequently, participants were required to rate their confidence and expertise using slider bars. The confidence slider had a range from “lower” on the left to “higher” on the right, while the expertise slider spanned from “not at all” on the left to “very much so” on the right, both internally implementing a 1-100 scaling. Additionally, participants indicated whether they had encountered the study previously before proceeding to the next trial. Upon completing the 11 trials, participants were debriefed on which trials they got correct and were subsequently asked to indicate whether they engaged in any form of cheating during the study. We hosted the study entirely on the Gorilla platform (?).
Exclusion criteria
Performance correlation between humans and LLMs
Fine-tuning on neuroscience corpora
To accomplish this, we employed the Low-Rank Adaptation of Large Language Models (LoRA) technique (Fig. S.7; (?)). LoRA efficiently extends the capabilities of general-purpose LLMs by introducing low-rank trainable parameters (referred to as “adapters”) into the existing model. This process effectively fine-tunes the model for downstream tasks without the need for prohibitively resource-intensive training of the entire model.
Training data
Training details
We fine-tuned Llama2-7b-chat from Meta (?) using weights available on Huggingface (https://huggingface.co/meta-llama/Llama-2-7b-chat-hf). We used a batch size of 1 and a chunk size of 2048. Training involved the use of the AdamW optimizer (?) with a learning rate of 2e-5 and gradient accumulation steps set at 8. A single training epoch was performed, along with a warm-up step of 0.03 and a weight decay rate of 0.001. The learning rate was controlled using a cosine learning rate scheduler. LoRA adapters, characterized by a rank of 8, an alpha value of 32, and a dropout rate of 0.1, were applied after all self-attention blocks and fully-connected layers. This results in total 17,891,328 trainable parameters, roughly $0.26\%$ of the entire parameters of the base model. To optimize training performance, bf16 mixed precision training and data parallelism were employed. We used 4 Nvidia A100 (80GB) GPUs hosted on the Microsoft Azure platform. An epoch of training takes roughly 42 GPU hours.
Evaluation
We tested the fine-tuned model on BrainBench using the same procedure as before. To verify the significance of performance improvement, we performed a paired t-test with respect to the difference in perplexity of incorrect and correct options before and after fine-tuning. We also performed a paired t-test with respect to the perplexity of the correct options before and after fine-tuning.
@@
Figure S.7: Low-Rank Adaptation (LoRA) for parameter-efficient large language model (LLM) fine-tuning. Image adapted from (?). Training LLMs from scratch can be computationally expensive, especially when the model has many billions of parameters and the training data are massive. Ideally, one would be able to take advantage of a previously trained LLM (i.e., a base model) and build from it. Toward this end, various Parameter Efficient Fine-Tuning (PEFT) techniques have been proposed. Rather than retrain the LLM, these techniques preserve the LLM’s original pre-trained weights and train a small subset of additional parameters that either enhance an existing capability in the LLM or introduce a new one. In our case, we tune a base LLM to the neuroscience literature. This tuning strategy significantly alleviates the computational burden while achieving comparable performance to training a new LLM from scratch. Among many PEFT approaches, we adopted LoRA in our study to fine-tune the 7-Billion parameter version of the LLaMa2 model on neuroscience abstracts. LoRA fixes the pretrained model weights and injects trainable rank decomposition matrices into each layer of the transformer. During training, the same input sequence (or intermediate layer output) $x$ is processed separately by the pretrained weights and the low-rank adaptation matrices ($A$ and $B$). The low-rank matrices possess the only trainable parameters in the model. The final output $h$ is computed as a coordinate-wise addition between the product of the pretrained weights and the adaptation matrices, which are further processed by subsequent layers.
@@
Tasks
Authors
Primary writing
Xiaoliang Luo, Bradley C. Love
Test case creation
Bati Yilmaz, Isil Poyraz Bilgin, Anton Pashkov
Tereza Okalova, Alexandra O. Cohen, Felipe Yáñez
Elkhan Yusifov, N Apurva Ratan Murty, Kangjoo Lee
Valentina Borghesani, Sepehr Razavi, Justin M. Ales
Rui Mata, Michael Gaebler, Guiomar Niso
Leyla Loued-Khenissi, Anna Behler, Kaustubh R. Patil
Mikail Khona, Roberta Rocca, Kevin K. Nejad
Alessandro Salatiello, Jonathan Nicholas
Daniele Marinazzo, Chloe M. Hall
Pui-Shee Lee, Sebastian Musslick
Nicholas E. Myers, Jennifer K Bizley, Sherry Dongqi Bao
Nianlong Gu, Jessica Dafflon, Bradley C. Love
Quality control
Bati Yilmaz, Kevin K. Nejad, Daniele Marinazzo
Pui-Shee Lee, Nianlong Gu, Chloe M. Hall
Kangjoo Lee, Sebastian Musslick, Akilles Rechardt
N Apurva Ratan Murty, Ilia Sucholutsky, Guiomar Niso
Isil Poyraz Bilgin, Rui Mata, Tereza Okalova
Mikail Khona, Justin M. Ales, Michael Gaebler
Elkhan Yusifov, Leyla Loued-Khenissi, Jessica Dafflon
Jonathan Nicholas, Felipe Yáñez, Roberta Rocca
Valentina Borghesani, Sherry Dongqi Bao
Alexandra O. Cohen, Alessandro Salatiello
Xiaoliang Luo, Bradley C. Love
GPT-4 case creation
Kevin K. Nejad, Xiaoliang Luo, Bradley C. Love
Human-machine teaming
Felipe Yáñez, Xiaoliang Luo, Bradley C. Love
LoRA fine tuning
Guangzhi Sun, Xiaoliang Luo, Bradley C. Love
Model evaluation
Xiaoliang Luo, Guangzhi Sun, Martin Ferianc
Bradley C. Love
Building the experiment
Akilles Rechardt, Xiaoliang Luo, Bradley C. Love
Data analysis
Xiaoliang Luo, Akilles Rechardt, Bradley C. Love
Conceptualization and strategy
Xiaoliang Luo, Bradley C. Love
Figure creation
Xiaoliang Luo, Bati Yilmaz, Isil Bilgin
Chloe M. Hall, Guiomar Niso, Bradley C. Love
Useful input and suggestions on project
All authors
Commenting and editing on the manuscript
All authors
Gaia Molinaro, Judy Zhu, Majd Abdallah, Yuri G. Pavlov, Jong-eun Lee, Adam Harris, Zhaoning Li, Roman Kessler, Lexi Zhang, Maciej Szul, Pranjul Gupta, Sunreeta Bhattacharya, Jellina Prinsen, Celine Gallagher, Michael Anes, Maarten Laroy, Tobias Ackels, Carina Forster, Pedro Gonçalves, Tommy Mcconnell, Diane Whitmer, Debottam Kundu, Benjamin Pasquereau, Jeremy Manning, Maciej Szul, Ahmed Hussain, Nicolas Clairis, Ignacio Vega-Vásquez, Kun Chen, Jeremy Hogeveen, Sina Salehi, Suseendrakumar Duraivel , Edgar Guevara, Ziyao Zhang, Thomas J. Younts, Marek Muszyński, Leonardo Dalla Porta, Todd Gureckis, Parnian Rafei, Feng-Chun Chou, Keith Temple, Alp Altunkaya, Andrew Tan, Jin Ho Yun, Arnau Marin-Llobet, Brian Lord, Daniel Lindh, Simon Besson-Girard, Eren Irmak, Emin Çelik, Aman Maharjan, Irene Sophia Plank
Table S.5: Participants acknowledgment. List of participants.

### v3

@@ Acknowledgments
This work was supported the ESRC (ES/W007347/1), Microsoft (Accelerate Foundation Models Research Program), and a Royal Society Wolfson Fellowship (18302) to B.C.L. We thank Mona Garvert, Pradeep Reddy Raamana, Todd Hare, Yoav Kessler, Oliver Robinson and D.R. for their assistance. We thank the participants of the online study: Gaia Molinaro, Judy Zhu, Majd Abdallah, Yuri G. Pavlov, Jong-eun Lee, Adam Harris, Zhaoning Li, Roman Kessler, Lexi Zhang, Maciej Szul, Pranjul Gupta, Sunreeta Bhattacharya, Jellina Prinsen, Celine Gallagher, Michael Anes, Maarten Laroy, Tobias Ackels, Carina Forster, Pedro Gonçalves, Tommy Mcconnell, Diane Whitmer, Debottam Kundu, Benjamin Pasquereau, Jeremy Manning, Maciej Szul, Ahmed Hussain, Nicolas Clairis, Ignacio Vega-Vásquez, Kun Chen, Jeremy Hogeveen, Sina Salehi, Suseendrakumar Duraivel , Edgar Guevara, Ziyao Zhang, Thomas J. Younts, Marek Muszyński, Leonardo Dalla Porta, Todd Gureckis, Parnian Rafei, Feng-Chun Chou, Keith Temple, Alp Altunkaya, Andrew Tan, Jin Ho Yun, Arnau Marin-Llobet, Brian Lord, Daniel Lindh, Simon Besson-Girard, Eren Irmak, Emin Çelik, Aman Maharjan, Irene Sophia Plank.
Declaration of interests
The authors declare no competing interests.
Author contributions
XL and BCL were responsible for primary writing. Test case creation was handled by BY, IPB, AP, TO, AOC, FY, EY, ARM, KL, VB, SR, JMA, RM, MG, GN, LLK, AB, KRP, MK, RR, KKN, AS, JN, DM, CMH, PSL, SM, NEM, JKB, SDB, NG, JD, and BCL. Quality control was conducted by BY, KKN, DM, PSL, NG, CMH, KL, SM, AR, ARM, IS, GN, IPB, RM, TO, MK, JMA, MG, EY, LLK, JD, JN, FY, RR, VB, SDB, AOC, AS, XL, and BCL. GPT-4 case creation was managed by KKN, XL, and BCL. Human-machine teaming involved FY, XL, and BCL. LoRA fine-tuning was performed by GS, XL, and BCL. Model evaluation was executed by XL, GS, MF, and BCL. Building the experiment was carried out by AR, XL, and BCL. Data analysis was done by XL, AR, and BCL. Conceptualization and strategy were undertaken by XL and BCL. Figure creation was the work of XL, BY, IB, CMH, GN, and BCL. Useful input and suggestions on the project were provided by all authors. Commenting and editing on the manuscript were also done by all authors. For a table breakdown, see Table S.5.
@@
We confirm that our research complies with all relevant ethical regulations; Experimental Psychology Ethics Board (University College London) approved the study protocol (ethics protocol: EP/2017/011); We confirm that informed consent was obtained from all human participants; Participant compensation is not applicable to the current study; None of our studies were preregistered.
@@
Co-authors (Table S.5) and GPT-4 (Azure OpenAI API; version 2023-05-15) created test cases that formed BrainBench. All test cases were sourced from Journal of Neuroscience abstracts published in 2023 under the Creative Commons Attribution 4.0 International License (CC-BY). The abstracts are organized into five sections, namely Behavioral/Cognitive, Systems/Circuits, Neurobiology of Disease, Development/Plasticity/Repair, and Cellular/Molecular. In constructing BrainBench, we incorporated a total of 200 test cases crafted by human experts and an additional 100 test cases generated by GPT-4 (Azure OpenAI API; version 2023-05-15). All test cases were subjected to extensive quality control by human experts and GPT-4. For the distribution of test cases among subfields, refer to Fig. 3 for human-created cases and Fig. S.22 for GPT-4 generated cases.
@@
We tested LLMs by adapting the Eleuther AI Language Model Evaluation Harness framework  [25], which evaluates LLMs using a multiple choice setting. We presented LLMs with two versions of the abstracts from each test case separately. We prefixed each abstract with the prompt “You are a neuroscientist with deep knowledge in neuroscience. Here is an abstract from a neuroscience publication:” and applied model-specific instruction templates where appropriate. We then measured the perplexity of both passages and used perplexity as the indicator of whether LLMs favor one abstract or the other.
@@
In addition, we tested models using abstracts whose results (in terms of complete sentences) are randomly swapped from abstracts within the same neuroscience subfield. Importantly, in these “swapped” abstracts, the number of results remained consistent with the original. We applied the swapping to both original and altered abstracts and re-evaluated LLMs’ performance.
@@
To address this concern, we employed a common method from the literature to determine whether a given text is part of LLM’s training data  [18, 27]. This method involves calculating the zlib entropy and the perplexity ratio (Eq. 3) of a text sequence to infer its membership status.
@@
Zlib entropy is computed using the zlib text compression algorithm  [28], which measures the level of uncertainty in a text when compressed. It’s a data-agnostic way of evaluating text. On the other hand, LLM perplexity depends on the specific training data and thus is data-dependent. In general, if a piece of text surprises zlib but not LLM, it’s likely part of the training data.
@@
Human Evaluation
Participants
Procedure
Subsequently, participants were required to rate their confidence and expertise using slider bars. The confidence slider had a range from “lower” on the left to “higher” on the right, while the expertise slider spanned from “not at all” on the left to “very much so” on the right, both internally implementing a 1-100 scaling. Additionally, participants indicated whether they had encountered the study previously before proceeding to the next trial. Upon completing the 11 trials, participants were debriefed on which trials they got correct and were subsequently asked to indicate whether they engaged in any form of cheating during the study. We hosted the study entirely on the Gorilla platform  [29].
Exclusion criteria
Performance correlation between humans and LLMs
Fine-tuning on neuroscience corpora
To accomplish this, we employed the Low-Rank Adaptation of Large Language Models (LoRA) technique (Fig. S.18;  [19]). LoRA efficiently extends the capabilities of general-purpose LLMs by introducing low-rank trainable parameters (referred to as “adapters”) into the existing model. This process effectively fine-tunes the model for downstream tasks without the need for prohibitively resource-intensive training of the entire model.
Training data
Training details
We fine-tuned Mistral-7B-v0.1  [30] using weights available on Huggingface (. We used a batch size of 1 and a chunk size of 2048. Training involved the use of the AdamW optimizer  [31] with a learning rate of 2e-5 and gradient accumulation steps set at 8. Two training epochs were performed, along with a warm-up step of 0.03 and a weight decay rate of 0.001. The learning rate was controlled using a cosine learning rate scheduler. LoRA adapters, characterized by a rank of 256, an alpha value of 512, and a dropout rate of 0.1, were applied after all self-attention blocks and fully-connected layers. This results in total 629,145,600 trainable parameters, roughly $8\%$ of the entire parameters of the base model. To optimize training performance, bf16 mixed precision training and data parallelism were employed. We used 4 Nvidia A100 (80GB) GPUs hosted on the Microsoft Azure platform. An epoch of training takes roughly 65 GPU hours.
Evaluation
We tested the fine-tuned model on BrainBench using the same procedure as before. To verify the significance of performance improvement, we performed a paired t-test with respect to the perplexity of the correct options before and after fine-tuning.
Relative perplexities evaluated at token and sentence levels
To gain further insights into how models solve BrainBench test cases, we broke down the change of relative perplexities over the sequence of tokens and sentences starting from the first alternation of results. Specifically, we looked at how often the model starts correctly at the first alternation and remains correct (Pos to Pos) or becomes incorrect (Pos to Neg) at the end. In addition, we categorized how often the model starts incorrectly but becomes correct (Neg to Pos) or remains incorrect (Neg to Neg).
Examples (Fig. S.9, S.10, S.11) show different patterns of relative perplexity changes between the altered and original abstracts over tokens. These examples demonstrate that the model does not always recognize the correct version at the first alternation, nor does it consistently remain in the correct (positive) or incorrect (negative) region across tokens.
To gain a holistic view, we examined all test cases and categorized how the model solves a question. Specifically, we looked at how often the model starts correctly at the first alternation and remains correct (Pos to Pos) or becomes incorrect (Pos to Neg) at the end. We also categorized how often the model starts incorrectly but becomes correct (Neg to Pos) or remains incorrect (Neg to Neg). The corresponding results are shown in Fig. S.15 (left panel). For more than half the abstracts, if the model is correct at the first alternation, it is likely to be correct at the end. Notably, $20\%$ of abstracts classified incorrectly at the first alternation are later classified correctly. We further evaluated beyond the first and final tokens by examining the number of sign flips of perplexity difference in between all tokens. As shown in Fig. S.15(right panel), the model exhibits fewer sign flips for abstracts eventually classified correctly than for those classified incorrectly. We performed the same analysis at the sentence level and observed similar patterns (Fig. S.12, S.13, S.14, S.16).
On the relationship between performance, model size and training data
What factors might be driving LLMs’ performance on BrainBench? As shown in Fig. 3A, models of varying sizes starting from 7 billion parameters perform similarly, with instruct/chat fine-tuned models performing worse and larger models do not necessarily achieve higher accuracy. To gain better insights, we trained a much smaller model, GPT-2 (124 million parameters) from scratch, entirely on the same neuroscience data utilized in this study. We trained dedicated tokenizer on the same data. We trained GPT-2 for 5 epochs until convergence and tested it on BrainBench. The model achieved $63\%$ accuracy  [32], which is very close to human-level performance but falling behind on bigger models. To map out the picture more clearly, we further trained GPT-2 medium (355M) and GPT-2 large (774M) following the same training procedure. They both outperformed GPT-2 (124M) with GPT-2 large achieving a $72\%$ accuracy further reducing the gap to much larger models we tested (see full results in Fig. S.2). This pattern of result suggests that the size of the model likely plays a critical role in integrating deep knowledge to provide accurate responses.
In addition, we tested another small LLM, TinyLLama (1.1B; v1.0), as a comparison, and the model achieved $70.5\%$ (base) and $65.5\%$ (chat) on BrainBench; falling much behind the 7B models we tested in the original study. Further, we conducted testing on Phi-3 (3.8B), a very recent (April, 2024) model developed by Microsoft renowned for its competitive edge even against models more than twice its size. Phi-3 achieved an $82.5\%$ accuracy on BrainBench. We attribute this performance to the known high-quality training data curated by Microsoft.
In light of all these results, we posit that achieving superhuman performance on BrainBench likely stems from a synergy between model size and training set quality. The Phi-3 result raises intriguing questions about the existence of a scaling law in this context. It prompts exploration into whether there’s a point of diminishing returns in model scaling and whether there exists a balance between model size, data quality and relevancy, and computational resources for optimal performance on BrainBench.
Tables
@@
RefinedWeb Known to be in Falcon’s training set  [33]
@@
Arxiv (Jun-Dec 2023) Known to not be in Llama2 and Galactica’s training set  [34, 20]
Biorxiv (Jun-Dec 2021) Known to be in Galactica’s training set  [20]
@@
Model/Human Coefficient Intercept Test Statistics
[…]
Table S.3: Logistic regression fits. For models, logistic regressions were fitted between perplexity differences of a test case and its correctness given a LLM. For human experts, logistic regressions were fitted between their confidence of a test case and its correctness. The significant positive correlations between model and human fits suggest that LLMs and humans are calibrated.
Tasks Authors
Primary writing Xiaoliang Luo, Bradley C. Love
Test case creation Bati Yilmaz, Isil Poyraz Bilgin, Anton Pashkov Tereza Okalova, Alexandra O. Cohen, Felipe Yáñez Elkhan Yusifov, N Apurva Ratan Murty, Kangjoo Lee Valentina Borghesani, Sepehr Razavi, Justin M. Ales Rui Mata, Michael Gaebler, Guiomar Niso Leyla Loued-Khenissi, Anna Behler, Kaustubh R. Patil Mikail Khona, Roberta Rocca, Kevin K. Nejad Alessandro Salatiello, Jonathan Nicholas Daniele Marinazzo, Chloe M. Hall Pui-Shee Lee, Sebastian Musslick Nicholas E. Myers, Jennifer K Bizley, Sherry Dongqi Bao Nianlong Gu, Jessica Dafflon, Bradley C. Love
Quality control Bati Yilmaz, Kevin K. Nejad, Daniele Marinazzo Pui-Shee Lee, Nianlong Gu, Chloe M. Hall Kangjoo Lee, Sebastian Musslick, Akilles Rechardt N Apurva Ratan Murty, Ilia Sucholutsky, Guiomar Niso Isil Poyraz Bilgin, Rui Mata, Tereza Okalova Mikail Khona, Justin M. Ales, Michael Gaebler Elkhan Yusifov, Leyla Loued-Khenissi, Jessica Dafflon Jonathan Nicholas, Felipe Yáñez, Roberta Rocca Valentina Borghesani, Sherry Dongqi Bao Alexandra O. Cohen, Alessandro Salatiello Xiaoliang Luo, Bradley C. Love
GPT-4 case creation Kevin K. Nejad, Xiaoliang Luo, Bradley C. Love
Human-machine teaming Felipe Yáñez, Xiaoliang Luo, Bradley C. Love
LoRA fine tuning Guangzhi Sun, Xiaoliang Luo, Bradley C. Love
Model evaluation Xiaoliang Luo, Guangzhi Sun, Martin Ferianc Bradley C. Love
Building the experiment Akilles Rechardt, Xiaoliang Luo, Bradley C. Love
Data analysis Xiaoliang Luo, Akilles Rechardt, Bradley C. Love
Conceptualization and strategy Xiaoliang Luo, Bradley C. Love
Figure creation Xiaoliang Luo, Bati Yilmaz, Isil Bilgin Chloe M. Hall, Guiomar Niso, Bradley C. Love
Useful input and suggestions on project All authors
Commenting and editing on the manuscript All authors
Figures
@@
Figure S.2: BrainBench performance across models of varying sizes and training data. Models with 7 billion parameters or more achieve similar results on BrainBench, while their instruct/chat fine-tuned variants perform systematically worse. The GPT-2 series (124M, 355M, 774M) trained entirely on the neuroscience literature from scratch shows progressively better results, matching or surpassing human performance and closing the gap to larger models. Phi3, with half the size of the 7B models, achieves competitive results likely due to its high-quality training data. In contrast, TinyLlama, with 1.1 billion parameters, lags behind on BrainBench. Overall, performance is influenced by both model size and the quality and relevance of the training data.
@@
Figure S.4: LLM performance with coherent, swapped and no context. There is a significant performance decline when evaluating with coherent versus swapped contexts, with further degradation when context is absent.
@@
Figure S.9: Perplexity difference at token level. Red dots indicate alternations. For this example, the model makes an incorrect response at the first alternation. As the model processes more tokens, the perplexity difference increases between the altered and original leading to a correct response at the end (Mistral-7B-v0.1).
Figure S.10: Perplexity difference at token level. Red dots indicate alternations. For this example, the model makes a correct response starting from the first alternation already and stays being correct throughout the rest of the tokens (Mistral-7B-v0.1).
Figure S.11: Perplexity difference at token level. Red dots indicate alternations. For this example, the model makes a correct response starting from the first alternation. The model switches to an incorrect response after having processed about half of the abstract. The model’s response flips once again at the final alternation and makes a correct response at the end (Mistral-7B-v0.1).
Figure S.12: Perplexity difference at sentence level. Red dots indicate the current sentence contains at least one alternations.
Figure S.13: Perplexity difference at sentence level. Red dots indicate the current sentence contains at least one alternations.
Figure S.14: Perplexity difference at sentence level. Red dots indicate the current sentence contains at least one alternations.
Figure S.15: Cumulative perplexity differences at token level. Left panel shows proportions of different decision flips of a model. A flip is determined by comparing the cumulative perplexity difference between the incorrect and the original abstracts until the first diverging token and the entire abstract. For example, a “Neg to Pos” flip suggests the model has a negative perplexity difference at the first diverging token and has a positive perplexity difference at the final token. Right panel shows the average number of perplexity difference sign flips in between tokens among abstracts classified correctly and incorrectly by the model.
Figure S.16: Cumulative perplexity differences at sentence level. Left panel shows proportions of different decision flips of a model. A flip is determined by comparing the cumulative perplexity difference between the incorrect and the original abstracts until the sentence that contains the first diverging token and the entire abstract. For example, a “Neg to Pos” flip suggests the model has a negative perplexity difference at the sentence that contains the first diverging token and has a positive perplexity difference at the final token. Right panel shows the average number of perplexity difference sign flips in between sentences among abstracts classified correctly and incorrectly by the model.
Figure S.17: Expertise distribution across subfields. The distribution of expertise are generally similar across subfields. The top 20% experts report expertise around or above 50/100.
@@
Figure S.18: Low-Rank Adaptation (LoRA) for parameter-efficient large language model (LLM) fine-tuning. Image adapted from  [19]. Training LLMs from scratch can be computationally expensive, especially when the model has many billions of parameters and the training data are massive. Ideally, one would be able to take advantage of a previously trained LLM (i.e., a base model) and build from it. Toward this end, various Parameter Efficient Fine-Tuning (PEFT) techniques have been proposed. Rather than retrain the LLM, these techniques preserve the LLM’s original pre-trained weights and train a small subset of additional parameters that either enhance an existing capability in the LLM or introduce a new one. In our case, we tune a base LLM to the neuroscience literature. This tuning strategy significantly alleviates the computational burden while achieving comparable performance to training a new LLM from scratch. Among many PEFT approaches, we adopted LoRA in our study to fine-tune the 7-Billion parameter version of the LLaMa2 model on neuroscience abstracts. LoRA fixes the pretrained model weights and injects trainable rank decomposition matrices into each layer of the transformer. During training, the same input sequence (or intermediate layer output) $x$ is processed separately by the pretrained weights and the low-rank adaptation matrices ($A$ and $B$). The low-rank matrices possess the only trainable parameters in the model. The final output $h$ is computed as a coordinate-wise addition between the product of the pretrained weights and the adaptation matrices, which are further processed by subsequent layers.